OpenStudy (anonymous):

Find the slope of the line through the pair of points. (-9,-4) (-9,8)

OpenStudy (experimentx):

(y2-y1)/(x2-x1) = 12/0 = infinity, thats ninety degree

OpenStudy (radar):

undefined slope, it is a vertical line